# P[$\arraybackslashp#1$

### v1

@@ 5.3 Further Discussion
We apply GhostSpec using Ghost-mse similarity to investigate the recently debated lineage of Pangu-Pro-MoE by comparing it with several open-source models, including the Qwen series and Llama variants.
@@
To examine whether spectral fingerprinting extends beyond attention layers, we analyze singular values of the MLP up-porj and down-proj weight matrices in each Transformer layer. We compute Ghost-mse similarity between Llama-2-7b-hf, its fine-tuned variants, and unrelated models.
As shown in Figure 6, MLP-based fingerprints effectively distinguish fine-tuned models (e.g., Vicuna-7b-v1.5, Llemma-7b) with similarity above 0.96 from unrelated models (e.g., Pythia-6.9b) showing near-zero similarity.
@@ 6 Conclusion
This paper tackles the challenge of verifying large language model lineage amid widespread undocumented reuse and potential plagiarism. We propose GhostSpec, a robust, data-independent, and efficient white-box method that constructs stable fingerprints from the singular value spectra of invariant attention matrix products, resilient to common model modifications and adversarial transformations. A penalty-based optimal path alignment algorithm enables handling architectural differences. Extensive experiments demonstrate that GhostSpec reliably identifies model ancestry across diverse transformations, including fine-tuning, pruning, merging, and architectural expansion. It accurately recovers known relationships and offers insights in complex cases. Moreover, the fingerprint’s tight coupling with model functionality makes evasion attempts impractical without degrading performance. Overall, GhostSpec offers a practical, trustworthy tool to protect intellectual property and improve transparency, advancing verifiable model provenance in open-source AI ecosystems.
@@ Appendix A Theoretical Foundation of the GhostSpec Fingerprint
This appendix provides the detailed theoretical underpinning for the robustness of the GhostSpec fingerprint, as claimed in Section 4.1 of the main paper. We first provide a rigorous derivation of the functionality-preserving transformations our method is designed to be invariant to. We then provide a formal proof of this invariance, taking into account the practical convention of storing transposed weight matrices.
@@
A.1 Functionality-Preserving Transformation Models
Our claim of robustness is based on the invariance of our fingerprint to transformations that modify model weights without altering the model’s input-output function. We now show in detail how these transformations preserve functionality.
FFN Permutation.
A standard FFN block with a gated linear unit (GLU) computes its output $y$ from an input $x$ as:
$y=\left(x\cdot W_{\text{up}}^{\top}\odot\sigma(x\cdot W_{\text{gate}}^{\top})\right)\cdot W_{\text{down}}^{\top},$
where $\odot$ is element-wise multiplication and $\sigma$ is an activation function.
An adversary can apply a permutation matrix $P$ to the intermediate neurons. The transformed (and stored) weights $\tilde{W}^{\top}$ are:
$\tilde{W}_{\text{up}}^{\top}=W_{\text{up}}^{\top}P,\quad\tilde{W}_{\text{gate}}^{\top}=W_{\text{gate}}^{\top}P,\quad\tilde{W}_{\text{down}}^{\top}=P^{-1}W_{\text{down}}^{\top}.$
The output with the transformed weights, $\tilde{y}$, is:
$\tilde{y}$ $=\left(x\cdot\tilde{W}_{\text{up}}^{\top}\odot\sigma(x\cdot\tilde{W}_{\text{gate}}^{\top})\right)\cdot\tilde{W}_{\text{down}}^{\top}$
$=\left(x\cdot W_{\text{up}}^{\top}P\odot\sigma(x\cdot W_{\text{gate}}^{\top}P)\right)\cdot(P^{-1}W_{\text{down}}^{\top})$
$=\left(\left(x\cdot W_{\text{up}}^{\top}\odot\sigma(x\cdot W_{\text{gate}}^{\top})\right)\cdot P\right)\cdot(P^{-1}W_{\text{down}}^{\top})$
$=\left(x\cdot W_{\text{up}}^{\top}\odot\sigma(x\cdot W_{\text{gate}}^{\top})\right)\cdot I\cdot W_{\text{down}}^{\top}=y.$
The output $y$ is identical, hence functionality is preserved.
@@ A.2 Formal Proof of Invariance
This section provides the formal proof for the invariance of our spectral fingerprints, which is a cornerstone of the method’s robustness. Our fingerprint is based on the singular value spectra of two specific matrix products, which we have selected to be strictly invariant to the corresponding transformations defined above.
@@
A.3 Compatibility with Multi-Head Attention and Positional Encodings
The functionality-preserving transformations and their corresponding invariance proofs described previously were presented on generic linear transformation layers. In modern Transformer architectures, these transformations must be compatible with key mechanisms such as Multi-Head Attention (MHA) and Rotary Positional Embeddings (RoPE). This section details how our method ensures this compatibility.
Multi-Head Attention (MHA).
In the multi-head attention mechanism, the Query, Key, and Value matrices ($Q$, $K$, $V$) are logically partitioned into $h$ heads, with each head operating independently in a subspace of the model’s hidden dimension. To apply transformations without disrupting this structure, our transformation matrices $C$ (for V-O) and $P$ (for Q-K) are constructed as block-diagonal matrices:
$C=\begin{pmatrix}C_{1}&0&\cdots&0\\ 0&C_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&C_{h}\end{pmatrix},\quad P=\begin{pmatrix}P_{1}&0&\cdots&0\\ 0&P_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&P_{h}\end{pmatrix}$
Here, each sub-matrix $C_{i}$ and $P_{i}$ has dimensions matching that of a single head (head_dim). This design has two critical advantages:
  1. 1.
    Isolated Transformations: Each sub-matrix operates solely within the subspace of its corresponding head, never mixing information across different heads. This guarantees that the independence of the multi-head mechanism is preserved.
  2. 2.
    Preservation of Invariance: Since the overall transformation is linear and the fundamental algebraic properties (e.g., $CC^{-1}=I$) hold for block-diagonal matrices, the proofs for functionality preservation and product invariance (e.g., $\widetilde{M}_{ov}=M_{ov}$) derived in the previous section remain fully valid for the entire weight matrix.
Rotary Positional Embedding (RoPE).
RoPE injects positional information by rotating pairs of dimensions within the Query and Key vectors. For instance, it rotates dimensions $(0,1)$, then independently rotates dimensions $(2,3)$, and so on. Any transformation applied to $W_{q}$ and $W_{k}$ must respect this pairwise structure to avoid corrupting the positional information.
To this end, we impose a stronger constraint on the Q-K transformation matrix $P$, making it doubly block-diagonal. Specifically, the transformation matrix for each head, $P_{i}$, is itself a block-diagonal matrix composed of $d_{\text{head}}/2$ smaller $2\times 2$ matrices:
$P_{i}=\begin{pmatrix}P_{i,1}^{(2\times 2)}&&0\\ &\ddots&\\ 0&&P_{i,d_{\text{head}}/2}^{(2\times 2)}\end{pmatrix}$
This meticulous structure ensures that our transformation operates only within the 2D subspaces on which RoPE acts, never mixing information between them (e.g., information from dimension 1 is not leaked to dimension 2). This guarantees that RoPE can still be applied correctly to the query and key vectors after our transformation, thus strictly preserving the model’s output.
Similarly, this more granular structural constraint does not alter the fact that $P$ is an invertible matrix. Therefore, the proof of invariance for our invariant product $M_{q^{\top}k}=W_{q}^{\top}W_{k}$ remains sound.
@@ B.1 Core Function Definitions
Distance correlation is used to compute the ‘GhostSpec-corr‘ similarity score (Section 4.2, Equation 7). It is a measure of statistical dependence between two random vectors of arbitrary, not necessarily equal, dimension. Unlike Pearson correlation, distance correlation is zero if and only if the vectors are statistically independent, making it a robust metric for comparing the complex, non-linear trends of our layer-wise spectral mean sequences.
@@ B.2 Hyperparameter Settings
The gap penalty $\rho$ for the Penalty-based Optimal Spectral Alignment (POSA) algorithm (Algorithm 1), which is central to comparing models of varying depth (Section 4.2), was set empirically. We observed that a value corresponding to the average single-layer MSE distance between moderately similar models provided a good balance between matching corresponding layers and allowing for necessary skips in pruned or expanded models. Across all experiments, a fixed value of $\rho=0.002$ was used.
@@
The parameters for the final inverted Sigmoid transformation (Equation 6), used to convert the raw path distance from POSA into the final ‘GhostSpec-mse‘ similarity score, were determined empirically. The goal was to select values that best separate the distance distributions of related and unrelated model pairs, thereby maximizing the F1-score on a validation set. Based on this analysis, we use a fixed threshold $\tau=0.00371$ and a steepness factor $k=1000$.
@@ Appendix C Detailed Experimental Setup
This appendix provides supplementary details for the experimental setup described in Section 5.1.
@@ C.1 Model Dataset
Our comprehensive dataset consists of 63 model pairs, created using **Llama-2-7b** and **Mistral-7B** as primary base models. This dataset covers a wide range of transformations, including fine-tuning, compression, merging, and expansion, with ground-truth labels indicating whether each pair is ’related’ or ’unrelated’. A complete list of all model pairs and their corresponding modification type is provided in Table 2.
Primary Model: Llama-2-7b Primary Model: Mistral-7B
vicuna-7b-v1.5 Fine-tuning Mistral-7B-Instruct-v0.1 Fine-tuning
llemma_7b Fine-tuning dolphin-2.2.1-mistral-7b Fine-tuning
CodeLlama-7b-hf Fine-tuning
chinese-llama-2-7b Fine-tuning
coma-7B-v0.1 Fine-tuning
Llama-2-chat-shortgpt-30 Structured Pruning
Llama-2-pruned50-retrained Unstructured Pruning Triunvirato-7b Merging & Expansion
Llama-2-pruned70-retrained Unstructured Pruning Daredevil-7B Merging & Expansion
GBLM-Pruner-LLaMA-2-7B Unstructured Pruning BioMistral-merged-instruct Merging & Expansion
llama2-7b-func-calling-slerp Merging & Expansion
FuseLLM-7B Merging & Expansion
LLaMA-Pro-8B Merging & Expansion
Camelidae-8x7B Merging & Expansion
Llama-2-7b-chat-scaled Adversarial Transforms Baichuan2-7B-Base Unrelated Models
Llama-2-7b-chat-permuted Adversarial Transforms internlm2-base-7b Unrelated Models
Mistral-7B-v0.1 Unrelated Models Qwen2.5-7B Unrelated Models
Qwen2.5-7B Unrelated Models pythia-6.9b Unrelated Models
Yi-1.5-6B Unrelated Models mpt-7b Unrelated Models
internlm2-base-7b Unrelated Models opt-6.7b Unrelated Models
opt-6.7b Unrelated Models Yi-1.5-6B Unrelated Models
deepseek-coder-6.7b Unrelated Models deepseek-coder-6.7b Unrelated Models
Qwen1.5-7B Unrelated Models Qwen1.5-7B Unrelated Models
@@ Appendix D Additional Experiments and Visualizations
D.1 Ablation Studies
@@
To demonstrate the critical role of our invariant product construction (Section 4.1), we performed an ablation study. We compared the performance of the full GhostSpec method against a “naive” variant that computes fingerprints directly from the singular values of the raw attention weights ($W_{q}$, $W_{k}$, $W_{v}$, $W_{o}$). The specific computational difference between the two methods is as follows:
@@
  • Naive Variant (Vulnerable Method): For the ablation study, we created a “naive” version that bypasses the invariant product construction. Instead, it directly computes the SVD of the four individual attention weight matrices for each layer $i$:
@@
We tested both methods on the task of identifying a transformed version of Llama-2-7b-chat-hf that had undergone both scaling and permutation attacks. The results are shown in Table LABEL:tab:ablation_results. The naive method fails completely, with similarity scores dropping significantly. In contrast, GhostSpec maintains near-perfect similarity, proving that the invariant product construction is essential for robustness against these attacks.
@@
We selected ‘Llama-2-7b‘ (32 layers) as the base model and compared it against a curated set of seven models, all of which have different layer counts, making a sophisticated alignment strategy essential. These models were categorized as:
  • Related Models: Five models derived from Llama-2 but with varying depths, including ‘LLaMA-Pro-8B‘ (40 layers), ‘Llama-2-7b-chat-hf-30-sparsity‘ (23 layers), and the ‘Sheared-LLaMA-1.3B‘ series (24 layers).
  • Unrelated Models: Two architecturally distinct models, ‘Qwen2-7B‘ (28 layers) and ‘Qwen2.5-7B‘ (28 layers).
@@
To quantitatively evaluate the discriminative power of each alignment strategy, we aggregated the results. For each method, we calculated the average similarity score for the “Related” group and the “Unrelated” group. The key metric for our evaluation is the Discriminative Gap ($\Delta$), which we define as the difference between these two average scores:
@@
A larger gap indicates a superior ability to clearly and reliably distinguish genuine derivatives from unrelated models. The results of this analysis are presented in Figure LABEL:fig:posa_ablation_correlation and Figure 7(b).
@@
D.2 Visualization of the Fingerprint Space

### v2

@@ 5.3 Further Discussion
We apply GhostSpec using GhostSpec-mse similarity to investigate the recently debated lineage of Pangu-Pro-MoE by comparing it with several open-source models, including the Qwen series and Llama variants.
@@
To examine whether spectral fingerprinting extends beyond attention layers, we analyze singular values of the MLP up-proj and down-proj weight matrices in each Transformer layer. We compute GhostSpec-mse similarity between Llama-2-7b-hf, its fine-tuned variants, and unrelated models.
As shown in Figure 6, MLP-based fingerprints effectively distinguish fine-tuned models with similarity above 0.96 from unrelated models showing near-zero similarity.
@@ 6 Conclusion
This paper tackles LLM lineage verification amid widespread reuse and potential plagiarism. We propose GhostSpec, a robust, data-independent, and efficient white-box method that constructs stable fingerprints from the singular value spectra of invariant attention matrix products. Resilient to common modifications and adversarial attacks, it employs a penalty-based optimal path alignment algorithm to handle architectural differences. Extensive experiments demonstrate GhostSpec reliably identifies model ancestry across diverse transformations, including fine-tuning, pruning, merging, and expansion. GhostSpec offers a practical, trustworthy tool to protect intellectual property and improve transparency in open-source AI ecosystems.
Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grant No. 62306216, the Fundamental Research Funds for the Central Universities under Grant No. 2042025kf0026, and the Technology Innovation Program of Hubei Province under Grant No. 2024BAB043.
@@ Appendix A Theoretical Foundation of the GhostSpec Fingerprint
We first characterize the class of potential functionality-preserving weight transformation attacks that can significantly alter model parameters without changing the input-output mapping. To counter these attacks, we identify specific matrix products that remain strictly unaffected. We then provide a formal proof of this invariance, taking into account the practical convention of storing transposed weight matrices.
@@
A.1 Functionality-Preserving Transformation Attack
@@ A.2 Formal Proof of Invariance
This section provides the formal proof that our spectral fingerprints remain invariant when subjected to the functionality-preserving transformation attacks defined above. This resistance is a cornerstone of the method’s robustness. We demonstrate that the singular value spectra of our selected matrix products are strictly immune to these adversarial manipulations.
@@ B.1 Core Function Definitions
Distance correlation is used to compute the GhostSpec-corr similarity score (Section 4.2, Equation 7). It is a measure of statistical dependence between two random vectors of arbitrary, not necessarily equal, dimension. Unlike Pearson correlation, distance correlation is zero if and only if the vectors are statistically independent, making it a robust metric for comparing the complex, non-linear trends of our layer-wise spectral mean sequences.
@@ B.2 Hyperparameter Settings
The gap penalty $\rho$ for the Penalty-based Optimal Spectral Alignment (POSA) algorithm (Algorithm 1), which is central to comparing models of varying depth, was set empirically. We observed that a value corresponding to the average single-layer MSE distance between moderately similar models provided a good balance between matching corresponding layers and allowing for necessary skips in pruned or expanded models. Across all experiments, a fixed value of $\rho=0.002$ was used.
@@
The parameters for the final inverted Sigmoid transformation (Equation 6), used to convert the raw path distance from POSA into the final GhostSpec-mse similarity score, were determined empirically. The goal was to select values that best separate the distance distributions of related and unrelated model pairs, thereby maximizing the F1-score on a validation set. Based on this analysis, we use a fixed threshold $\tau=0.00371$ and a steepness factor $k=1000$.
@@ Appendix C Detailed Experimental Setup
This appendix provides supplementary details for the experimental setup described in Section 5.
@@ C.1 Model Dataset
Our comprehensive dataset is constructed using Llama-2-7B and Mistral-7B as the primary base models. This dataset covers a wide range of transformations, including fine-tuning, compression, merging, and expansion, with ground-truth labels indicating whether each pair is related or unrelated. A complete list of all model pairs and their corresponding modification type is provided in Table 2.
@@ C.2 Computing Infrastructure
C.3 Computational Efficiency Analysis
We analyze the computational efficiency of GhostSpec in two distinct stages: the one-time fingerprint extraction and the pairwise similarity comparison.
Theoretical Complexity.
  • Stage 1: Fingerprint Extraction. For a model with $L$ layers and hidden dimension $d$, we compute two invariant products per layer ($M_{qk}$ and $M_{vo}$) and perform Singular Value Decomposition (SVD). While the worst-case complexity for full SVD is $O(d^{3})$, our practical implementation employs randomized truncated SVD to extract only the top-$K$ singular values, where $K\ll d$ (typically determined by effective rank). This reduces the practical complexity to $O(L\cdot d^{2}K)$. Importantly, this is a one-time cost per model checkpoint.
  • Stage 2: Similarity Comparison. To compare two models with $N$ and $M$ layers, the POSA algorithm computes a pairwise distance matrix and executes dynamic programming to find the optimal alignment. The complexity is $O(N\cdot M\cdot K)$. Crucially, this step is independent of the model’s large hidden dimension $d$, making the retrieval and verification process extremely fast.
Empirical Runtime.
We measured execution times on a server equipped with a single NVIDIA GeForce RTX 4090 GPU (24GB VRAM) for extraction and standard Intel Xeon CPUs for comparison. Table 3 details the fingerprint extraction times for models of varying sizes.
Model Parameters Extraction Time (s)
Qwen3-1.7B 1.7B 30.7
Qwen3-4B 4B 46.6
Qwen3-8B 8B 132.0
Table 3: One-time fingerprint extraction times measured on a single RTX 4090 GPU. The process is efficient, taking approximately 2 minutes for an 8B parameter model.
The pairwise comparison stage is near-instantaneous. For example, computing the similarity between Llama-2-7B and Qwen2.5-7B takes approximately 0.24 seconds, and comparing Llama-2-7B with the significantly deeper Qwen2.5-32B takes only 0.62 seconds. This efficiency demonstrates that GhostSpec is highly scalable for checking a query model against a large repository of existing fingerprints.
@@ Appendix D Additional Experiments and Visualizations
D.1 Out-of-Sample Threshold Generalization
To evaluate the generalizability of our method and the robustness of the optimal thresholds derived in Section 5.1 ($\tau_{\text{mse}}=0.85$ and $\tau_{\text{corr}}=0.61$), we conducted out-of-sample evaluations on two additional model families with significantly different parameter scales: the Qwen2.5-3B family and the Qwen2.5-32B family.
We applied the fixed thresholds determined solely from the initial Llama-2/Mistral-7B dataset.
Results on 3B Scale.
Table 4 presents the results for the 3B parameter scale. GhostSpec correctly identifies all fine-tuned variants as related, with scores well above the thresholds. Conversely, unrelated models like MiniMA-3B and bloom-3b are correctly rejected.
Base Model Comparison Model GhostSpec-mse GhostSpec-corr Ground Truth
Qwen2.5-3B Qwen2.5-3B-Instruct 0.9761 0.9999 Related
Qwen2.5-3B Qwen2.5-Coder-3B 0.9618 0.9407 Related
Qwen2.5-3B SDLM-3B-D4 0.9760 0.9990 Related
Qwen2.5-3B calme-3.1-instruct-3b 0.9761 0.9999 Related
Qwen2.5-3B MiniMA-3B 0.0000 0.6097 Unrelated
Qwen2.5-3B bloom-3b 0.0000 0.2765 Unrelated
Table 4: Out-of-sample evaluation on the Qwen2.5-3B family using fixed thresholds ($\tau_{\text{mse}}>0.85$, $\tau_{\text{corr}}>0.61$). Scores in bold indicate a positive classification (Related) consistent with Ground Truth.
Results on 32B Scale.
Table 5 extends this validation to the larger 32B parameter scale. Despite the significant increase in model depth and width compared to the 7B base models used for calibration, GhostSpec maintains its discriminative power. All derivative models are identified with high confidence (MSE >0.97>0.97), while unrelated large models show negligible similarity.
Base Model Comparison Model GhostSpec-mse GhostSpec-corr Ground Truth
Qwen2.5-32B BFS-Prover-V2-32B 0.9761 0.9999 Related
Qwen2.5-32B K2-Think 0.9761 0.9996 Related
Qwen2.5-32B Qwen2.5-32B-Instruct 0.9761 0.9999 Related
Qwen2.5-32B SDLM-32B-D4 0.9761 0.9965 Related
Qwen2.5-32B Yi-34B 0.0003 0.2999 Unrelated
Qwen2.5-32B opus-v1-34B 0.0003 0.2934 Unrelated
Table 5: Out-of-sample evaluation on the Qwen2.5-32B family using fixed thresholds ($\tau_{\text{mse}}>0.85$, $\tau_{\text{corr}}>0.61$). The method remains robust at larger scales.
These results confirm that the structural fingerprints captured by GhostSpec are scale-invariant and that the decision thresholds established on 7B models generalize effectively to both smaller (3B) and larger (32B) architectures without the need for recalibration.
D.2 Ablation Studies
@@
To demonstrate the critical role of our invariant product construction (Section 4.1), we performed an ablation study. We compared the performance of the full GhostSpec method against a naive variant that computes fingerprints directly from the singular values of the raw attention weights ($W_{q}$, $W_{k}$, $W_{v}$, $W_{o}$). The specific computational difference between the two methods is as follows:
@@
  • Naive Variant (Vulnerable Method): For the ablation study, we created a naive version that bypasses the invariant product construction. Instead, it directly computes the SVD of the four individual attention weight matrices for each layer $i$:
@@
We tested both methods on the task of identifying a transformed version of Llama-2-7b-chat-hf that had undergone both scaling and permutation attacks. The results are shown in Table 6. The naive method fails completely, with similarity scores dropping significantly. In contrast, GhostSpec maintains near-perfect similarity, proving that the invariant product construction is essential for robustness against these attacks.
@@
Ablation on Invariant Product Components
To justify the design choice of using two invariant products ($M_{qk}=W_{q}W_{k}^{\top}$ and $M_{vo}=W_{v}W_{o}$) rather than a single one, we performed an ablation study on 28 model pairs (20 related, 8 unrelated). We evaluated the discriminative power of fingerprints derived from:
  1. 1.
    QK Product Only: Using only the singular values of $M_{qk}$.
  2. 2.
    VO Product Only: Using only the singular values of $M_{vo}$.
  3. 3.
    Both (Proposed): Using the concatenated fingerprints from both products.
As shown in Table 7, while each product independently captures meaningful structural information, combining them yields the highest discriminative gap for both metrics. This confirms that the $Q$-$K$ and $V$-$O$ interactions encode complementary information about the attention mechanism’s identity, and using both provides the most robust fingerprint.
Fingerprint Configuration $\Delta_{\text{mse}}$ $\Delta_{\text{corr}}$
QK Product Only ($M_{qk}$) 0.7099 0.3962
VO Product Only ($M_{vo}$) 0.5739 0.4568
Both (Proposed) 0.7370 0.4672
Table 7: Ablation study on the contribution of invariant matrix products. Combining both QK and VO products consistently yields the highest discriminative gap ($\Delta$).
@@
We selected Llama-2-7b (32 layers) as the base model and compared it against a curated set of seven models, all of which have different layer counts, making a sophisticated alignment strategy essential. These models were categorized as:
  • Related Models: Five models derived from Llama-2 but with varying depths, including LLaMA-Pro-8B (40 layers), Llama-2-7b-chat-hf-30-sparsity (23 layers), and the Sheared-LLaMA-1.3B series (24 layers).
  • Unrelated Models: Two architecturally distinct models, Qwen2-7B (28 layers) and Qwen2.5-7B (28 layers).
@@
To quantitatively evaluate the discriminative power of each alignment strategy, we aggregated the results. For each method, we calculated the average similarity score for the Related group and the Unrelated group. The key metric for our evaluation is the Discriminative Gap ($\Delta$), which we define as the difference between these two average scores:
@@
A larger gap indicates a superior ability to clearly and reliably distinguish genuine derivatives from unrelated models. The results of this analysis are presented in Figure 7(a) and Figure 7(b).
@@
Sensitivity Analysis of POSA Penalty ($\rho$)
The Penalty-based Optimal Spectral Alignment (POSA) algorithm relies on a gap penalty parameter, $\rho$, to balance the cost of matching dissimilar layers against the cost of skipping layers (introducing gaps). To validate our choice of $\rho=0.002$ and analyze the method’s sensitivity to this hyperparameter, we conducted an experiment using the identical curated dataset of 7 model pairs described in the ablation study above. This dataset consists of the base Llama-2-7b model compared against the same group of 5 related variants with differing depths and 2 architecturally distinct unrelated models.
We measured the Discriminative Gap ($\Delta$), defined as the difference between the average similarity score of related models and unrelated models:
$\Delta=\text{AvgScore}(\text{Related})-\text{AvgScore}(\text{Unrelated}).$ (22)
A higher $\Delta$ indicates better separation capability.
Table 8 presents the results for $\rho\in[0.0,0.01]$. The results reveal a stable high-performance plateau for $\rho\in[0.0,0.002]$, where the discriminative gap remains maximal. Beyond this range (e.g., $\rho\geq 0.005$), the penalty becomes too severe, forcing suboptimal alignments and reducing the gap, particularly for the MSE-based metric. Our default choice of $\rho=0.002$ sits at the edge of this stable region, providing an optimal balance between alignment flexibility and structural constraint.
Gap Penalty ($\rho$) $\Delta_{\text{mse}}$ $\Delta_{\text{corr}}$ Observation
0.000 0.9206 0.6385 Stable Plateau
0.001 0.9194 0.6385 Stable Plateau
0.002 (Default) 0.9194 0.6386 Optimal Choice
0.005 0.8447 0.6381 Onset of Decline
0.010 0.8074 0.6340 Significant Decline
Table 8: Sensitivity analysis of the POSA gap penalty $\rho$. The Discriminative Gap ($\Delta$) remains stable for small penalties but degrades as $\rho$ increases, justifying our selection of 0.002.
D.3 Visualization of the Fingerprint Space